# COLT: Towards Completeness-Oriented Tool Retrieval for Large Language Models

### v1

@@ 6. Conclusion
Appendix
Appendix A Details of ToolLens Construction
The prompts designed for GPT-4 to construct ToolLens are shown in Table 7.
Appendix B Case Study
B.1. Comparison of ToolLens and ToolBench.
Scene | ToolLens | ToolBench |
Cooking | I’m planning a meal using the ingredient beef. | I’m organizing a dinner party for my friends and I need some recipe suggestions. Please provide me with a variety of chicken recipes and their nutritional information. Also, I would appreciate some cocktail recommendations to complement the meal. |
Travel | I want to travel to Paris. | I’m planning a family vacation to London in August. Can you help me find the best hotels in London for a family of four? Also, provide me with the distance between Birmingham, Alabama, and Sacramento, California. Lastly, recommend some fun activities for kids in London. |
Investment | I am currently tracking my cryptocurrency investments. | I’m a teacher and I want to plan an engaging lesson on current events for my students. Can you provide me with specific articles related to bitcoin from different news sources? Additionally, I’d like to gather information about crypto news and the latest trends in the market. Finally, could you recommend any climate change news from reliable sources to discuss the impact on the economy? |
Table 6 displays a comparison of three typical instances of user queries in the ToolLens and ToolBench datasets. The queries within ToolLens, crafted through our innovative five-step methodology, exhibit notable distinctions from those in the ToolBench dataset, underscoring the superiority of our dataset. As can be observed, the queries within ToolLens are concise, appearing more natural and multifaceted. For example, in the travel scene, the query from ToolLens may simply express a desire to visit Paris, making it more natural and multifaceted. Conversely, the ToolBench dataset contains queries that explicitly seek information about hotels, distances and tourist attractions.
B.2. Comparison of Responses
Table 8 displays the comparison of responses generated by GPT-4, which integrates information from multiple tools retrieved through different tool retrievers in response to user queries. As can be observed, due to the more comprehensive and complete tools provided by COLT, the responses generated by GPT-4 are more comprehensive and effective.
Appendix C Complete Response
The complete responses illustrated in the introduction are detailed in Table 9. We can find that the absence of any essential tool from the ground-truth tool set compromises the ability to fully address user queries. Moreover, the transition from having two missing tools to having two incorrect ones results in a marked decline in the quality of responses generated by GPT-4. This observation demonstrates that providing redundant tools does not necessarily facilitate problem-solving. Such shortcomings impede successful query resolution, emphasizing the necessity of having both complete and accurate tools to achieve optimal performance.
I: Scene Mining |
/* I: Task prompt */ Your task is to generate potential scenes where a specific tool function might be used. Below is the description of the tool function. Please provide a scene where this tool function could be utilized. Just give the scene, no explanation needed. Here is an example: /* Example */ The given tool description is: {category_name:Food, tool_name:Worldwide Recipes, api_name: Suggestions, api_description: Get Suggestions, required_parameters:[ {"name": "q", "type": "STRING", "description": "", "default": "chicken" }], optional_parameters: [], method: GET, template_response: {"success": "bool", "message": "str", "results": {"ingredients": ["list of str with length 413"], "searches": ["list of str with length 10"]}}} Your response would be: cooking. /* Generation of Scene */ The API description is: … Your response is: |
II: Query Generation (Single-Tool) |
/* Task prompt */ Your task is to create a potentially multifaceted user query based on a given scene. The query should be presented as a declarative sentence. It should include both required and optional parameters, for which you need to generate random information. For required parameters like IP address, location, coordinates, etc., provide specific details. For example, instead of simply stating “an address”, provide the exact road and district names. Here is an example: /* Example */ The given scene is Going out fishing. The given required and optional parameters are as follows: required_parameters: [{“name”: “longitude”,“type”: “STRING”,“description”: “”,“default”: “-79.99”},{“name”: “date”, “type”: “STRING”,“description”: “”,“default”: “2022-01-01”}, {“name”:“latitude”,“type”: “STRING”,“description”: “”,“default”: “39.53”}], “optional_parameters”: [] Your response would be: I’m planning a fishing trip at the coordinates 51.5074 N, -0.1278 W for the date 2020-06-15. /* Generation of Query */ The scenario for this task is: … The required and optional parameters are as follows: … Your response is: |
III: Tool Aggregation (Category Generation) |
/* Task prompt */ You will be provided with a user query, and your task is to provide three types of API functions that can better answer this query. /* Generation of Categories */ The user query is: … Your response is: |
IV: Query Rewritting (Multi-Tool) |
/* Task prompt */ Your task is to rewrite the user query based on the given user query and the required parameters. The original user query may not include all the required parameters, and you need to change it to a user query that includes all the required parameters, for which you need to generate random information. For required parameters like IP address, location, coordinates, etc., provide specific details. For example, instead of simply stating ‘an address’, provide the exact road and district names. Try to rewrite the user query as if it were a human inquiry, without dividing it into three separate parts. /* Generation of Query */ The user query is: … The required parameters are as follows: … Your response is: |
Query |
Input query: I’m planning a picnic outing in Helsinki. |
BM25: |
That sounds wonderful! Helsinki has a number of beautiful parks that would be perfect for a picnic. Just to name a few, there are the Esplanade Park, Kaisaniemi Park, and Kaivopuisto Park. These parks have areas with picnic tables and open fields where you could spread out a blanket. Enjoy your picnic! |
ANCE: |
Planning a picnic in Helsinki can be a wonderful experience, as the city offers many beautiful parks and natural settings. Here are some steps and tips to help you plan your outing:1. Choose a Date and Check the Weather: Helsinki’s weather can be quite variable, so check the forecast to choose a date with pleasant weather. During summer, the days are warm and long, making it an ideal time for picnics. 2. Select a Location: Some popular picnic spots in Helsinki include: Suomenlinna: A sea fortress and UNESCO World Heritage site offering green areas with sea views. Kaivopuisto: A park with wide green spaces, views of the sea, and nice spots for picnics. Enjoy your picnic in Helsinki, and remember to soak in the beauty of the Finnish capital! |
TAS-B: |
That sounds like a great plan! Here are some wonderful picnic spots in Helsinki: 1. Kaivopuisto Park: This park is one of the oldest and best known in Helsinki, providing stunning views over the sea. 2. Esplanade Park: Located in the city center with lots of cafes and shops nearby. Concerts and events often take place here during summer. 3. Suomenlinna: A picnic on the fortress island of Suomenlinna is a unique experience. After a picnic, you can tour the historic fortifications and visit the museums on the island. 4. Seurasaari Island: A great place for an outdoor picnic. You can also tour the open-air museum and feed the squirrels on this island. 5. Hietaniemi Beach: If you want a picnic by the beach, this is a great spot. It’s a great place for swimming, too. Enjoy your picnic! |
coCondensor: |
When planning a picnic outing in Helsinki, consider the following tips: 1. Check the Weather: Helsinki has a northern continental or subarctic climate, so always check the weather forecast ahead of time. Summers can be pleasant for picnics, usually from late May through August, with temperatures ranging from 15°C to 25°C (59°F to 77°F). Remember that weather can be unpredictable, so it’s good to have a backup plan in case of rain. 2. Select a Location: Some popular parks and coastal areas in Helsinki perfect for picnics are: Kaivopuisto: Located by the sea, offering lovely views and open green spaces. Suomenlinna: A sea fortress that’s been turned into a park, accessible by ferry, providing a historic setting for a picnic. Töölönlahti: A bay area surrounded by a park in the center of the city, ideal for a relaxed outdoor meal. Enjoy your picnic outing, and remember to be prepared for any changes in weather, protect the environment, and savor the local flavors! |
Contriever: |
Planning a picnic outing in Helsinki is a great idea, especially during the warmer months. Here’s a helpful guide to assist you in organizing a lovely picnic: 1. Check the Weather: Make sure you monitor the weather forecast for Helsinki, so you pick a day with pleasant weather. The best months for a picnic are typically May through September when the temperature is moderate and the risk of rain is lower. 2. Select a Location: Kaivopuisto Park: Offers beautiful sea views and lots of green spaces for a picnic. Suomenlinna: A sea fortress island that provides historical sights as well as picnic areas. Töölönlahti Bay: Surrounded by a park that features walking paths and green areas perfect for picnicking. Esplanadi Park: Located in the heart of Helsinki, this park is perfect for people-watching and enjoying urban nature. Seurasaari Island: Gives you a taste of Finnish nature and traditional architecture, with plentiful picnic spots. 3. Prepare Your Picnic Basket: Fill it with snacks and food items that do not require immediate refrigeration. Don’t forget to bring drinks, and if you fancy something local, pick up some Finnish rye bread and cold cuts for sandwiches, or grab Karelian pastries from a local bakery. Enjoy your picnic! |
COLT: |
When planning a picnic outing in Helsinki, here are some steps and tips to consider: 1. Choose a Location: Helsinki offers many stunning spots for picnic outings. Some of the popular locations include: Suomenlinna Island: A UNESCO World Heritage site accessible by ferry, offering beautiful views and green spaces. Kaivopuisto: A seaside park with large grassy areas, perfect for a picnic and views of the sea. Central Park (Keskuspuisto): A large park area offering a natural setting for a quiet and relaxing picnic. Töölönlahti Bay: A bay area surrounded by parks where you can enjoy a waterside picnic. 2. Check the Weather: The weather in Helsinki can be unpredictable, so before you head out, check the current weather conditions and forecast. Make sure you bring appropriate clothing for the weather, and consider bringing an umbrella or waterproof blanket if there’s a chance of rain. 3. Pack Your Picnic: Prepare a basket or cooler with your favorite picnic foods. Don’t forget to pack: Food: Sandwiches, fruits, salads, snacks, etc. Beverages: Water, juice, or soft drinks; and if you desire, a bottle of wine or champagne (but make sure to check local laws regarding alcohol consumption in public spaces). Utensils: Plates, cutlery, napkins, and cups if needed. Extras: Condiments, a cutting board, a knife, and a bottle opener. 4. Bring Comfort Items: For a comfortable picnic, consider packing: A large blanket to sit on. Cushions or foldable chairs for extra comfort. Sunscreen and sunglasses on a sunny day. 5. Entertainment and Games: Change things up from just eating, to also having some fun. Bring along: A frisbee, soccer ball, or other sports equipment. A kite if it’s a windy day. Board games, playing cards, or books for a relaxed time. 6. Respect the Environment: When you’re ready to leave, make sure to clean up after yourself. Bring trash bags and recycle any waste if possible. 7. Local Amenities: Check if there are nearby restrooms and if pets are allowed if you plan to bring one. 8. Transportation: Helsinki has an excellent public transportation system including buses, trams, metro, and ferries. If you plan to visit somewhere like Suomenlinna, check the ferry schedule ahead of time. For locations within the city, you can use HSL’s journey planner to find the best route. Now you’re all set for your picnic in Helsinki! Enjoy the beautiful scenery and the tasty food with good company. |
Query |
Input query: I would like to know the value of 5 ounces of gold plus 1 million AMZN stocks in CNY. |
With Complete Tools: |
First, let’s calculate the value of 5 ounces of gold in USD: 5 ounces * 1919.7605807522993 USD/ounce = 9598.802903761497 USD. Then, let’s calculate the value of 1 million AMZN stocks in USD: 1,000,000 stocks * 130 USD/stock = 130,000,000 USD. Now, let’s convert these values to YNC using the exchange rate: 9598.802903761497 USD * 7.24 YNC/USD = 69,495.48654289563 YNC for the gold. 130,000,000 USD * 7.24 YNC/USD = 941,200,000 YNC for the stocks. So, the value of 5 ounces of gold is approximately 69,495.49 YNC and the value of 1 million AMZN stocks is approximately 941,200,000 YNC. |
One Is Incorrect: |
Sorry, I can’t provide the current value as I am an AI model trained by OpenAI and I don’t have real-time data or internet access to look up current stock prices or gold prices. However, you can calculate it by finding the current price of gold per ounce and the current stock price of AMZN, then multiply by the amount you have and the current exchange rate. |
Two Are Incorrect: |
Sorry, I am an AI model developed by OpenAI and I don’t have real-time data or the ability to provide current stock prices or the value of gold. I recommend checking a reliable financial news website or a stock market app for the most accurate and up-to-date information. |
One Is Missing: |
The value of 5 ounces of gold in YNC would be calculated by first finding the value in USD and then converting it to YNC.5 ounces of gold in USD = 5 * XAU USD = 5 * 1919.7605807522993 = 9598.802903761497 USD. Then, convert this to YNC using the exchange rate: 9598.802903761497 USD * 7.24 YNC/USD = 69,495.4893405 YNC. As for the value of 1 million AMZN stocks, we would need the current price of AMZN stocks in USD to calculate it. Without that information, we cannot provide the value in YNC. |
Two Are Missing: |
As an AI, I don’t have real-time data or future predictions. However, I can tell you how to calculate the value based on the information you provided. First, calculate the value of 5 ounces of gold: 5 ounces * 1919.7605807522993 USD/ounce = 9598.802903761497 USD. Then, calculate the value of 1 million AMZN stocks: 1,000,000 stocks * 130 USD/stock = 130,000,000 USD. Finally, to convert these values to YNC, you would need the exchange rate between USD and YNC. Unfortunately, I don’t have that information. Once you have the exchange rate, you can multiply the values in USD by the exchange rate to get the values in YNC. |
Without Tools: |
Sorry, but as an AI, I’m not able to provide real-time data or future predictions for stock prices or the value of gold. I recommend checking a reliable financial news source or financial services website for the most current information. |

### v2

@@ 6. Conclusion
Acknowledgements.
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No. 62377044), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, funds for building world-class universities (disciplines) of Renmin University of China, and PCC@RUC.References
- (1)
- Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774 (2023).
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems 33 (2020), 1877–1901.
- Chiang et al. (2023) Wei-Lin Chiang, Zhuohan Li, Zi Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E. Gonzalez, Ion Stoica, and Eric P. Xing. 2023. Vicuna: An Open-Source Chatbot Impressing GPT-4 with 90%* ChatGPT Quality. https://vicuna.lmsys.org
- Chowdhery et al. (2023) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. 2023. Palm: Scaling language modeling with pathways. Journal of Machine Learning Research 24, 240 (2023), 1–113.
- Dai et al. (2023) Sunhao Dai, Ninglu Shao, Haiyuan Zhao, Weijie Yu, Zihua Si, Chen Xu, Zhongxiang Sun, Xiao Zhang, and Jun Xu. 2023. Uncovering ChatGPT’s Capabilities in Recommender Systems. In Proceedings of the 17th ACM Conference on Recommender Systems (RecSys ’23). ACM. https://doi.org/10.1145/3604915.3610646
- Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023. QLoRA: Efficient Finetuning of Quantized LLMs. arXiv:2305.14314 [cs.LG]
- Gao and Callan (2021) Luyu Gao and Jamie Callan. 2021. Unsupervised corpus aware language model pre-training for dense passage retrieval. arXiv preprint arXiv:2108.05540 (2021).
- Gao et al. (2024) Shen Gao, Zhengliang Shi, Minghang Zhu, Bowen Fang, Xin Xin, Pengjie Ren, Zhumin Chen, Jun Ma, and Zhaochun Ren. 2024. Confucius: Iterative Tool Learning from Introspection Feedback by Easy-to-Difficult Curriculum. In AAAI.
- Guo et al. (2022) Jiafeng Guo, Yinqiong Cai, Yixing Fan, Fei Sun, Ruqing Zhang, and Xueqi Cheng. 2022. Semantic models for the first-stage retrieval: A comprehensive review. ACM Transactions on Information Systems (TOIS) 40, 4 (2022), 1–42.
- Gutmann and Hyvärinen (2010) Michael Gutmann and Aapo Hyvärinen. 2010. Noise-contrastive estimation: A new estimation principle for unnormalized statistical models. In Proceedings of the thirteenth international conference on artificial intelligence and statistics. JMLR Workshop and Conference Proceedings, 297–304.
- He et al. (2020) Xiangnan He, Kuan Deng, Xiang Wang, Yan Li, Yongdong Zhang, and Meng Wang. 2020. Lightgcn: Simplifying and powering graph convolution network for recommendation. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval. 639–648.
- Hofstätter et al. (2021) Sebastian Hofstätter, Sheng-Chieh Lin, Jheng-Hong Yang, Jimmy Lin, and Allan Hanbury. 2021. Efficiently teaching an effective dense retriever with balanced topic aware sampling. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval. 113–122.
- Hou et al. (2024) Yupeng Hou, Junjie Zhang, Zihan Lin, Hongyu Lu, Ruobing Xie, Julian McAuley, and Wayne Xin Zhao. 2024. Large Language Models are Zero-Shot Rankers for Recommender Systems. arXiv:2305.08845 [cs.IR]
- Huang et al. (2023) Yue Huang, Jiawen Shi, Yuan Li, Chenrui Fan, Siyuan Wu, Qihui Zhang, Yixin Liu, Pan Zhou, Yao Wan, Neil Zhenqiang Gong, and Lichao Sun. 2023. MetaTool Benchmark: Deciding Whether to Use Tools and Which to Use. arXiv preprint arXiv: 2310.03128 (2023).
- Izacard et al. (2021) Gautier Izacard, Mathilde Caron, Lucas Hosseini, Sebastian Riedel, Piotr Bojanowski, Armand Joulin, and Edouard Grave. 2021. Unsupervised dense information retrieval with contrastive learning. arXiv preprint arXiv:2112.09118 (2021).
- Järvelin and Kekäläinen (2002) Kalervo Järvelin and Jaana Kekäläinen. 2002. Cumulated gain-based evaluation of IR techniques. ACM Transactions on Information Systems (TOIS) 20, 4 (2002), 422–446.
- Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova. 2019. Bert: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of naacL-HLT, Vol. 1. 2.
- Li et al. (2023) Minghao Li, Yingxiu Zhao, Bowen Yu, Feifan Song, Hangyu Li, Haiyang Yu, Zhoujun Li, Fei Huang, and Yongbin Li. 2023. API-Bank: A Comprehensive Benchmark for Tool-Augmented LLMs. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 3102–3116. https://doi.org/10.18653/v1/2023.emnlp-main.187
- Liang et al. (2023) Percy Liang, Rishi Bommasani, Tony Lee, Dimitris Tsipras, Dilara Soylu, Michihiro Yasunaga, Yian Zhang, Deepak Narayanan, Yuhuai Wu, Ananya Kumar, Benjamin Newman, Binhang Yuan, Bobby Yan, Ce Zhang, Christian Cosgrove, Christopher D. Manning, Christopher Ré, Diana Acosta-Navas, Drew A. Hudson, Eric Zelikman, Esin Durmus, Faisal Ladhak, Frieda Rong, Hongyu Ren, Huaxiu Yao, Jue Wang, Keshav Santhanam, Laurel Orr, Lucia Zheng, Mert Yuksekgonul, Mirac Suzgun, Nathan Kim, Neel Guha, Niladri Chatterji, Omar Khattab, Peter Henderson, Qian Huang, Ryan Chi, Sang Michael Xie, Shibani Santurkar, Surya Ganguli, Tatsunori Hashimoto, Thomas Icard, Tianyi Zhang, Vishrav Chaudhary, William Wang, Xuechen Li, Yifan Mai, Yuhui Zhang, and Yuta Koreeda. 2023. Holistic Evaluation of Language Models. arXiv:2211.09110 [cs.CL]
- Liu et al. (2023b) Nelson F. Liu, Kevin Lin, John Hewitt, Ashwin Paranjape, Michele Bevilacqua, Fabio Petroni, and Percy Liang. 2023b. Lost in the Middle: How Language Models Use Long Contexts. arXiv:2307.03172 [cs.CL]
- Liu et al. (2023a) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu. 2023a. G-Eval: NLG Evaluation using GPT-4 with Better Human Alignment. arXiv:2303.16634 [cs.CL]
- Lu et al. (2022) Yao Lu, Max Bartolo, Alastair Moore, Sebastian Riedel, and Pontus Stenetorp. 2022. Fantastically Ordered Prompts and Where to Find Them: Overcoming Few-Shot Prompt Order Sensitivity. arXiv:2104.08786 [cs.CL]
- Mallen et al. (2022) Alex Mallen, Akari Asai, Victor Zhong, Rajarshi Das, Daniel Khashabi, and Hannaneh Hajishirzi. 2022. When not to trust language models: Investigating effectiveness of parametric and non-parametric memories. arXiv preprint arXiv:2212.10511 (2022).
- Mialon et al. (2023) Grégoire Mialon, Roberto Dessì, Maria Lomeli, Christoforos Nalmpantis, Ram Pasunuru, Roberta Raileanu, Baptiste Rozière, Timo Schick, Jane Dwivedi-Yu, Asli Celikyilmaz, et al. 2023. Augmented language models: a survey. arXiv preprint arXiv:2302.07842 (2023).
- Paranjape et al. (2023) Bhargavi Paranjape, Scott Lundberg, Sameer Singh, Hannaneh Hajishirzi, Luke Zettlemoyer, and Marco Tulio Ribeiro. 2023. Art: Automatic multi-step reasoning and tool-use for large language models. arXiv preprint arXiv:2303.09014 (2023).
- Parisi et al. (2022) Aaron Parisi, Yao Zhao, and Noah Fiedel. 2022. Talm: Tool augmented language models. arXiv preprint arXiv:2205.12255 (2022).
- Qin et al. (2023a) Yujia Qin, Shengding Hu, Yankai Lin, Weize Chen, Ning Ding, Ganqu Cui, Zheni Zeng, Yufei Huang, Chaojun Xiao, Chi Han, et al. 2023a. Tool learning with foundation models. arXiv preprint arXiv:2304.08354 (2023).
- Qin et al. (2023b) Yujia Qin, Shihao Liang, Yining Ye, Kunlun Zhu, Lan Yan, Yaxi Lu, Yankai Lin, Xin Cong, Xiangru Tang, Bill Qian, et al. 2023b. Toolllm: Facilitating large language models to master 16000+ real-world apis. arXiv preprint arXiv:2307.16789 (2023).
- Qu et al. (2024) Changle Qu, Sunhao Dai, Xiaochi Wei, Hengyi Cai, Shuaiqiang Wang, Dawei Yin, Jun Xu, and Ji-Rong Wen. 2024. Tool Learning with Large Language Models: A Survey. arXiv preprint arXiv:2405.17935 (2024).
- Robertson et al. (2009) Stephen Robertson, Hugo Zaragoza, et al. 2009. The probabilistic relevance framework: BM25 and beyond. Foundations and Trends® in Information Retrieval 3, 4 (2009), 333–389.
- Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language models can teach themselves to use tools. arXiv preprint arXiv:2302.04761 (2023).
- Shen et al. (2024) Yongliang Shen, Kaitao Song, Xu Tan, Dongsheng Li, Weiming Lu, and Yueting Zhuang. 2024. Hugginggpt: Solving ai tasks with chatgpt and its friends in hugging face. Advances in Neural Information Processing Systems 36 (2024).
- Song et al. (2023) Yifan Song, Weimin Xiong, Dawei Zhu, Cheng Li, Ke Wang, Ye Tian, and Sujian Li. 2023. Restgpt: Connecting large language models with real-world applications via restful apis. arXiv preprint arXiv:2306.06624 (2023).
- Sottana et al. (2023) Andrea Sottana, Bin Liang, Kai Zou, and Zheng Yuan. 2023. Evaluation Metrics in the Era of GPT-4: Reliably Evaluating Large Language Models on Sequence to Sequence Tasks. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 8776–8788. https://doi.org/10.18653/v1/2023.emnlp-main.543
- Sparck Jones (1972) Karen Sparck Jones. 1972. A statistical interpretation of term specificity and its application in retrieval. Journal of documentation 28, 1 (1972), 11–21.
- Sun et al. (2023) Weiwei Sun, Zheng Chen, Xinyu Ma, Lingyong Yan, Shuaiqiang Wang, Pengjie Ren, Zhumin Chen, Dawei Yin, and Zhaochun Ren. 2023. Instruction Distillation Makes Large Language Models Efficient Zero-shot Rankers. arXiv:2311.01555 [cs.IR]
- Tang et al. (2024) Jiakai Tang, Sunhao Dai, Zexu Sun, Xu Chen, Jun Xu, Wenhui Yu, Lantao Hu, Peng Jiang, and Han Li. 2024. Towards Robust Recommendation via Decision Boundary-aware Graph Contrastive Learning. Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (2024).
- Tang et al. (2023a) Qiaoyu Tang, Ziliang Deng, Hongyu Lin, Xianpei Han, Qiao Liang, and Le Sun. 2023a. ToolAlpaca: Generalized Tool Learning for Language Models with 3000 Simulated Cases. arXiv preprint arXiv:2306.05301 (2023).
- Tang et al. (2023b) Raphael Tang, Xinyu Zhang, Xueguang Ma, Jimmy Lin, and Ferhan Ture. 2023b. Found in the Middle: Permutation Self-Consistency Improves Listwise Ranking in Large Language Models. arXiv:2310.07712 [cs.CL]
- Thakur et al. (2021) Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna Gurevych. 2021. BEIR: A Heterogeneous Benchmark for Zero-shot Evaluation of Information Retrieval Models. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2). https://openreview.net/forum?id=wCu6T5xFjeJ
- Touvron et al. (2023) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288 (2023).
- Vu et al. (2023) Tu Vu, Mohit Iyyer, Xuezhi Wang, Noah Constant, Jerry Wei, Jason Wei, Chris Tar, Yun-Hsuan Sung, Denny Zhou, Quoc Le, et al. 2023. Freshllms: Refreshing large language models with search engine augmentation. arXiv preprint arXiv:2310.03214 (2023).
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35 (2022), 24824–24837.
- Wu and Aji (2023) Minghao Wu and Alham Fikri Aji. 2023. Style Over Substance: Evaluation Biases for Large Language Models. arXiv:2307.03025 [cs.CL]
- Xiong et al. (2020) Lee Xiong, Chenyan Xiong, Ye Li, Kwok-Fung Tang, Jialin Liu, Paul Bennett, Junaid Ahmed, and Arnold Overwijk. 2020. Approximate nearest neighbor negative contrastive learning for dense text retrieval. arXiv preprint arXiv:2007.00808 (2020).
- Xu et al. (2023) Qiantong Xu, Fenglu Hong, Bo Li, Changran Hu, Zhengyu Chen, and Jian Zhang. 2023. On the Tool Manipulation Capability of Open-source Large Language Models. arXiv preprint arXiv:2305.16504 (2023).
- Yao et al. (2022) Shunyu Yao, Jeffrey Zhao, Dian Yu, Nan Du, Izhak Shafran, Karthik Narasimhan, and Yuan Cao. 2022. React: Synergizing reasoning and acting in language models. arXiv preprint arXiv:2210.03629 (2022).
- Ye et al. (2024) Junjie Ye, Guanyu Li, Songyang Gao, Caishuang Huang, Yilong Wu, Sixian Li, Xiaoran Fan, Shihan Dou, Qi Zhang, Tao Gui, et al. 2024. Tooleyes: Fine-grained evaluation for tool learning capabilities of large language models in real-world scenarios. arXiv preprint arXiv:2401.00741 (2024).
- Yuan et al. (2024) Siyu Yuan, Kaitao Song, Jiangjie Chen, Xu Tan, Yongliang Shen, Ren Kan, Dongsheng Li, and Deqing Yang. 2024. EASYTOOL: Enhancing LLM-based Agents with Concise Tool Instruction. arXiv preprint arXiv:2401.06201 (2024).
- Zhao et al. (2023) Wayne Xin Zhao, Jing Liu, Ruiyang Ren, and Ji-Rong Wen. 2023. Dense Text Retrieval based on Pretrained Language Models: A Survey. ACM Trans. Inf. Syst. (dec 2023).
- Zheng et al. (2024) Yuanhang Zheng, Peng Li, Wei Liu, Yang Liu, Jian Luan, and Bin Wang. 2024. ToolRerank: Adaptive and Hierarchy-Aware Reranking for Tool Retrieval. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING) (2024).
- Zhu (2004) Mu Zhu. 2004. Recall, precision and average precision. Department of Statistics and Actuarial Science, University of Waterloo, Waterloo 2, 30 (2004), 6.